A gas undergoes adiabatic transformation using the first law of thermodynamics show that this process is represented by the equation pVy =constant
1
Expert's answer
2020-08-17T09:46:34-0400

The first law of thermodynamics says

"\delta Q=dU+pdV"

For adiabatic processes

"\delta Q=0"

Hence

"dU+pdV=0"

"\left(\frac{\partial U}{\partial T}\right)_VdT+\frac{\left[\left(\frac{\partial U}{\partial V}\right)_TdT+p\right]}{\left(\frac{\partial T}{\partial V}\right)_p}dV=0""C_VdT+\frac{\left[C_p-C_V\right]}{\left(\frac{\partial T}{\partial V}\right)_p}dV=0"

For perfect gas


"pV=RT"

"C_VdT+\frac{\left[C_p-C_V\right]}{\frac{V}{T}}dV=0"

"\frac{dT}{T}+\left[C_p/C_V-1\right]\frac{dV}{V}=0"

"\frac{dT}{T}+\left[\gamma-1\right]\frac{dV}{V}=0"

After integrating

"\ln T+(\gamma-1)\ln V=\ln \rm Const"

"TV^{\gamma-1}=\rm Const"

or


"\frac{pV}{R}V^{\gamma-1}=\rm Const"

Finally


"pV^{\gamma}=\rm Const"